Solve each system by addition:
x+y=7
x-y=9

Answer by checkley77(12844)

x+y=7
x-y=9
-----------
2x=16
x=16/2
x=8 ans.
8+y=7
y=7-8
y=-1 ans.
Proof:
8-(-1)=9
8+1=9
9=9